# Topic Modeling as Multi-Objective Contrastive Optimization

### v2

@@ 1 Introduction
shuttle lands on planet
job career ask development
0.0093 0.0026
star astronaut planet light moon
0.8895 0.9178
shuttle lands on planet zeppelin/scardino
job career ask development zeppelin/scardino
0.9741/0.9413 0.0064/0.0080
star astronaut planet light moon
0.1584/0.2547 0.8268/0.7188
@@ 3.3 Gradient-based Multi-Objective Optimization for Multi-Task Learning
[…]
@@ 5.2 Main Results
Table 4: Document classification results with topic representations. Method 20NG IMDb AG News BATM 30.8 66.0 56.0 SCHOLAR 32.2 83.4 74.8 SCHOLAR + BAT 52.9 73.1 59.0 NTM+CL 54.4 84.2 76.2 HyperMiner 43.0 71.3 63.4 WeTe 48.5 80.3 65.7 Our Model 57.0 86.4 78.6
Table 5: Ablation Results on the Objective Control Strategy. Method Wiki $T=50$ $T=200$ UW 0.482$\pm$0.009 0.474$\pm$0.005 GradNorm 0.485$\pm$0.007 0.469$\pm$0.006 PCGrad 0.489$\pm$0.011 0.472$\pm$0.005 RW 0.492$\pm$0.009 0.476$\pm$0.007 Linear-$\alpha=0.25$ 0.483$\pm$0.011 0.487$\pm$0.007 Linear-$\alpha=0.50$ 0.491$\pm$0.015 0.485$\pm$0.004 Our Model 0.496$\pm$0.010 0.491$\pm$0.006
@@
Document classification. We evaluate the effectiveness of our topic representations for the downstream task. Particularly, we extract latent distributions inferred by NTMs in the $T=50$ setting and train Random Forest classifiers with the number of decision trees of 100 to predict the category of each document. Because the Wiki dataset does not include a corresponding label to each input document, we use 20NG, IMDb, and AGNews for this experiment. As shown in Table 5.2, our model accomplishes the best performance over other NTMs, substantiating the refined usefulness of the topic representations yielded by our model.
@@ 5.3 Ablation Study
Objective control strategy. We perform an ablation study and analyze the importance of our proposed multi-objective training framework. In detail, we remove the Pareto stationary solver and then conduct the training with the approaches of linear combination, in which we attach the weight $\alpha$ to the contrastive objective with $\alpha\in\{0.25,0.5\}$, the Uncertainty Weighting (UW) (Kendall et al., 2018), gradient normalization (GradNorm) (Chen et al., 2018), Projecting Conflicting Gradients (PCGrad) (Yu et al., 2020), and Random Weighting (RW) (Lin et al., 2021). Without the multi-objective framework, the contrastive NTM suffers from sub-optimal performance, as shown in Table 5.2. Hypothetically, this means that in the fixed linear and heuristics weighting systems, the optimization procedure could not efficiently adapt to the phenomenon that the contrastive objective overwhelms the topic modeling task. Hence, contrasting samples might supply excessive mutual information which eclipses the useful features for topic learning. Due to the length limit, more experiments upon the objective control strategy can be referred to Appendix B.
@@
MinPooling
0.332$\pm$0.003 0.280$\pm$0.005 0.190$\pm$0.005 0.186$\pm$0.004 0.491$\pm$0.007 0.481$\pm$0.004 0.318$\pm$0.015 0.210$\pm$0.002
MeanPooling
0.334$\pm$0.005 0.283$\pm$0.005 0.194$\pm$0.006 0.195$\pm$0.010 0.492$\pm$0.006 0.484$\pm$0.009 0.321$\pm$0.014 0.211$\pm$0.002
SumPooling
0.333$\pm$0.004 0.281$\pm$0.003 0.191$\pm$0.007 0.188$\pm$0.005 0.493$\pm$0.013 0.484$\pm$0.004 0.322$\pm$0.006 0.215$\pm$0.003
GAP
0.339$\pm$0.004 0.288$\pm$0.004 0.197$\pm$0.003 0.196$\pm$0.005 0.495$\pm$0.009 0.485$\pm$0.007 0.323$\pm$0.007 0.221$\pm$0.003
FSP
0.339$\pm$0.007 0.286$\pm$0.003 0.199$\pm$0.004 0.195$\pm$0.007 0.494$\pm$0.012 0.484$\pm$0.007 0.324$\pm$0.012 0.221$\pm$0.003
PSWE
0.334$\pm$0.006 0.284$\pm$0.004 0.194$\pm$0.006 0.191$\pm$0.003 0.494$\pm$0.013 0.484$\pm$0.009 0.323$\pm$0.006 0.218$\pm$0.001
@@
Pooling functions. We experiment with different pooling operations for negative samples $\varphi^{-}$ and compare their performance with our choice of MaxPooling. The experiment of pooling functions for positive samples $\varphi^{+}$ can be found in Appendix C. We adapt MeanPooling, SumPooling, Global Attention Pooling (GAP) (Li et al., 2016), Featurewise Sort Pooling (FSP) (Zhang et al., 2018), and Pooling by Sliced-Wasserstein Embedding (PSWE) (Naderializadeh et al., 2021). As can be observed from Table 6, MaxPooling acquires the highest topic coherence quality. We hypothesize that different from other aggregation functions, MaxPooling directly retrieves strong features, thus being able to extract important information and suppress the noise. We also realize that our proposed framework consistently outperforms the individual-based contrastive NTM, demonstrating the robustness of setwise contrastive learning with a variety of aggregation functions.
@@ Appendix C Further Study of Pooling Function for Positive Set Representation
In this appendix, we extend our ablation study of different choices of pooling functions $\varphi^{-}$ for negative set in Section 5.3 to pooling function $\varphi^{+}$ for positive set representation. In particular, similar to Section 5.3, we also compare our choice of pooling function for positive sets, i.e. MinPooling, with MaxPooling, MeanPooling, SumPooling, GAP (Li et al., 2016), FSP (Zhang et al., 2018), and PSWE (Naderializadeh et al., 2021). The results are denoted in Table 10.
@@
MinPooling
0.333$\pm$0.006 0.284$\pm$0.003 0.195$\pm$0.006 0.191$\pm$0.010 0.494$\pm$0.007 0.484$\pm$0.007 0.324$\pm$0.009 0.211$\pm$0.001
MeanPooling
0.335$\pm$0.003 0.284$\pm$0.005 0.197$\pm$0.006 0.194$\pm$0.007 0.492$\pm$0.009 0.481$\pm$0.006 0.324$\pm$0.013 0.213$\pm$0.002
SumPooling
0.336$\pm$0.006 0.286$\pm$0.004 0.195$\pm$0.007 0.186$\pm$0.008 0.493$\pm$0.009 0.481$\pm$0.008 0.320$\pm$0.013 0.212$\pm$0.001
GAP
0.336$\pm$0.006 0.289$\pm$0.005 0.196$\pm$0.006 0.186$\pm$0.005 0.492$\pm$0.010 0.482$\pm$0.007 0.318$\pm$0.009 0.210$\pm$0.002
FSP
0.336$\pm$0.004 0.285$\pm$0.003 0.197$\pm$0.006 0.191$\pm$0.010 0.494$\pm$0.012 0.484$\pm$0.008 0.323$\pm$0.014 0.218$\pm$0.002
PSWE
0.333$\pm$0.007 0.285$\pm$0.005 0.198$\pm$0.005 0.187$\pm$0.006 0.494$\pm$0.009 0.484$\pm$0.005 0.322$\pm$0.008 0.211$\pm$0.002
@@ Appendix E Examples of data augmentation methods
ronaldo poaches the points as owen tastes first real.
ronaldo’s scoring prowess was on full display as he secured the victory for real, while owen experienced his first taste of success with the team.
the serene sunset over the ocean painted the sky with hues of orange and pink, casting a peaceful aura over the coastal town.
i quite often find that when a film doesn’t evolve around a famous actor or actress but rather a story or style.
focusing on a compelling story or unique cinematic style can often lead to a more engaging movie experience, regardless of the star power of the actors involved.
the cat jumped over the fence and landed gracefully in the neighbor’s yard.
finally, a way to Google your hard drive You can usually find information more easily on the vast internet.
with this new tool, you’ll be able to search your hard drive just as efficiently as you browse the vast expanse of the internet.
the sound of laughter echoed through the empty park.
bioshield effort is inadequate, a study says experts say the bioshield program will not be adequate to create new vaccines and drugs to protect the nation against either a biological attack or natural epidemic.
the study’s findings cast doubt on the effectiveness of the bioshield program in developing the necessary vaccines and drugs to safeguard the nation from biological threats, according to experts.
the chef carefully placed the garnish on the plate, completing the presentation of the dish.
@@
ronaldo poaches the points as owen tastes first real.
cristiano cambone this rebounds one bellamy distinctly fourth gives.
maradona lem following 18 own ritchie mixes 1998 rather.
i quite often find that when a film doesn’t evolve around a famous actor or actress but rather a story or style.
never seemed ones way has he with screenplay reason we alter apart with notable film rather girlfriend only meant with explains rather simple.
get rarely perhaps still actually turn put episode come nothing conversely instead put besides scriptwriter there ms. while instance put inspired there quirky.
finally, a way to Google your hard drive You can usually find information more easily on the vast internet.
once both with it move online get turning stopping maybe if instead way account still quickly which this wealth websites.
turned instead put once rest lets want lot hard lot these less still link longer none last following boasts msn.
bioshield effort is inadequate, a study says experts say the bioshield program will not be adequate to create new vaccines and drugs to protect the nation against either a biological attack or natural epidemic.
nobelity helping part sufficiently both with suggests explained researchers know this nobelity planning ready any even providing move rather its doses made treatment move threatened this entire losing however with possessed fire rather unique malaria.
circumvents promises once poorly instead put applied nothing impact n’t following circumvents expanded step thought cannot suitable rest core introduced medication few antiviral rest need following strongest twice as put analyzing enemy there hence spreads.
@@
ronaldo poaches the points as owen tastes first real.
ronaldinho strafaci which point also keane taste second kind.
artd newsfeatures yahd nzse40 404-526-5456 3,399 leel arabised highjump.
i quite often find that when a film doesn’t evolve around a famous actor or actress but rather a story or style.
me very sometimes how but then an movie if know evolving where an famed starred instead actor though merely an book instead styles.
statistique kunzig -96 deylam suhn micale lacques nzse40 xinglong myoo fdch lacques -5000 basse-normandie zahr debehogne cetacea lacques arabised zahr tih.
finally, a way to google your hard drive you can usually find information more easily on the vast internet.
again . an so take yahoo my turn drives ’ll able typically how source than either . which surrounding web.
sahk lacques -96 yahd raht 1-732-390-4480 milanka musicnotes.com 1-732-390-4480 photoexpress suhn deylam tahn 100.00 zhahn romanosptimes.com newsfeatures puh condita.
bioshield effort is inadequate, a study says experts say the bioshield program will not be adequate to create new vaccines and drugs to protect the nation against either a biological attack or natural epidemic.
toshka efforts this insufficient . an studies thinks specialists believe which toshka programs would be not sufficient take creating for vaccine well drug take protecting which country opponents instead an chemical attacks instead abundant outbreak.
chivenor wmorrisglobe.com nehn europe/africa sahk lacques ortsgemeinden telegram.com gahl e-bangla newsfeatures chivenor sholdersptimes.com ortsgemeinden milanka 47-42-80-44 sylviidae yahd troo yanow chron.com bizmags tilove yahd telegram.com newsfeatures juh speech/language tvcolumn lacques fassihi 2378 zahr berris bse-100.
@@ Appendix F Statistical Significance Tests towards Topic Quality
In this appendix, to more rigorously assess the quality of our generated topics compared with baseline models, we conduct a statistical t-test for performance comparison in Section 5. In particular, we select the second-best method in all Tables, i.e. NTM+CL, Linear-$\alpha$ = 0.5, NTM+CL, NTM+CL, GAP, and GAP in Table 5.2, 5.2, 2, 3, 6, and 10, respectively. We provide our obtained $p$-values in Table 15, 16, 17, 18, 19, and 20. The results show that our improvement is statistically significant with all $p$-value $<0.05$.

### v3

@@ 1 Introduction
shuttle lands on planet job career ask development 0.0093 0.0026
star astronaut planet light moon 0.8895 0.9178
shuttle lands on planet zeppelin/scardino job career ask development zeppelin/scardino 0.9741/0.9413 0.0064/0.0080
star astronaut planet light moon 0.1584/0.2547 0.8268/0.7188
@@ 3.3 Gradient-based Multi-Objective Optimization for Multi-Task Learning
(Sener & Koltun, 2018) Let $\theta_{\text{shared}}^{*}$ denote a Pareto point. Then, there exists non-negative scalars $\{\alpha_{m}\}_{m=1}^{M},m\in\{1,2,…,M\}$ such that :
$\sum_{m=1}^{M}\alpha_{m}=1\quad\text{and}\quad\sum_{m=1}^{M}\alpha_{m}\nabla_{\theta_{\text{shared}}}\mathcal{L}_{m}(\theta_{\text{shared}}^{*},\theta^{i})=0.$ (4)
@@ 5.2 Main Results
Table 4: Document classification results with topic representations.
Method 20NG IMDb AG News
BATM 30.8 66.0 56.0
SCHOLAR 32.2 83.4 74.8
SCHOLAR + BAT 52.9 73.1 59.0
NTM+CL 54.4 84.2 76.2
HyperMiner 43.0 71.3 63.4
WeTe 48.5 80.3 65.7
Our Model 57.0 86.4 78.6
Table 5: Ablation Results on the Objective Control Strategy.
Method Wiki
$T=50$ $T=200$
UW 0.482$\pm$0.009 0.474$\pm$0.005
GradNorm 0.485$\pm$0.007 0.469$\pm$0.006
PCGrad 0.489$\pm$0.011 0.472$\pm$0.005
RW 0.492$\pm$0.009 0.476$\pm$0.007
Linear-$\alpha=0.25$ 0.483$\pm$0.011 0.487$\pm$0.007
Linear-$\alpha=0.50$ 0.491$\pm$0.015 0.485$\pm$0.004
Our Model 0.496$\pm$0.010 0.491$\pm$0.006
@@
Document classification. We evaluate the effectiveness of our topic representations for the downstream task. Particularly, we extract latent distributions inferred by NTMs in the $T=50$ setting and train Random Forest classifiers with the number of decision trees of 100 to predict the category of each document. Because the Wiki dataset does not include a corresponding label to each input document, we use 20NG, IMDb, and AGNews for this experiment. As shown in Table LABEL:tab:exp_dc, our model accomplishes the best performance over other NTMs, substantiating the refined usefulness of the topic representations yielded by our model.
@@ 5.3 Ablation Study
Objective control strategy. We perform an ablation study and analyze the importance of our proposed multi-objective training framework. In detail, we remove the Pareto stationary solver and then conduct the training with the approaches of linear combination, in which we attach the weight $\alpha$ to the contrastive objective with $\alpha\in\{0.25,0.5\}$, the Uncertainty Weighting (UW) (Kendall et al., 2018), gradient normalization (GradNorm) (Chen et al., 2018), Projecting Conflicting Gradients (PCGrad) (Yu et al., 2020), and Random Weighting (RW) (Lin et al., 2021). Without the multi-objective framework, the contrastive NTM suffers from sub-optimal performance, as shown in Table LABEL:tab:exp_ablation_objective_control_strategy. Hypothetically, this means that in the fixed linear and heuristics weighting systems, the optimization procedure could not efficiently adapt to the phenomenon that the contrastive objective overwhelms the topic modeling task. Hence, contrasting samples might supply excessive mutual information which eclipses the useful features for topic learning. Due to the length limit, more experiments upon the objective control strategy can be referred to Appendix B.
@@
MinPooling 0.332$\pm$0.003 0.280$\pm$0.005 0.190$\pm$0.005 0.186$\pm$0.004 0.491$\pm$0.007 0.481$\pm$0.004 0.318$\pm$0.015 0.210$\pm$0.002
MeanPooling 0.334$\pm$0.005 0.283$\pm$0.005 0.194$\pm$0.006 0.195$\pm$0.010 0.492$\pm$0.006 0.484$\pm$0.009 0.321$\pm$0.014 0.211$\pm$0.002
SumPooling 0.333$\pm$0.004 0.281$\pm$0.003 0.191$\pm$0.007 0.188$\pm$0.005 0.493$\pm$0.013 0.484$\pm$0.004 0.322$\pm$0.006 0.215$\pm$0.003
GAP 0.339$\pm$0.004 0.288$\pm$0.004 0.197$\pm$0.003 0.196$\pm$0.005 0.495$\pm$0.009 0.485$\pm$0.007 0.323$\pm$0.007 0.221$\pm$0.003
FSP 0.339$\pm$0.007 0.286$\pm$0.003 0.199$\pm$0.004 0.195$\pm$0.007 0.494$\pm$0.012 0.484$\pm$0.007 0.324$\pm$0.012 0.221$\pm$0.003
PSWE 0.334$\pm$0.006 0.284$\pm$0.004 0.194$\pm$0.006 0.191$\pm$0.003 0.494$\pm$0.013 0.484$\pm$0.009 0.323$\pm$0.006 0.218$\pm$0.001
@@
Pooling functions. We experiment with different pooling operations for negative samples $\varphi^{-}$ and compare their performance with our choice of MaxPooling. The experiment of pooling functions for positive samples $\varphi^{+}$ can be found in Appendix C. We adapt MeanPooling, SumPooling, Global Attention Pooling (GAP) (Li et al., 2016), Featurewise Sort Pooling (FSP) (Zhang et al., 2018), and Pooling by Sliced-Wasserstein Embedding (PSWE) (Naderializadeh et al., 2021). As can be observed from Table LABEL:tab:exp_pooling_results_negative, MaxPooling acquires the highest topic coherence quality. We hypothesize that different from other aggregation functions, MaxPooling directly retrieves strong features, thus being able to extract important information and suppress the noise. We also realize that our proposed framework consistently outperforms the individual-based contrastive NTM, demonstrating the robustness of setwise contrastive learning with a variety of aggregation functions.
@@ Appendix C Further Study of Pooling Function for Positive Set Representation
In this appendix, we extend our ablation study of different choices of pooling functions $\varphi^{-}$ for negative set in Section 5.3 to pooling function $\varphi^{+}$ for positive set representation. In particular, similar to Section 5.3, we also compare our choice of pooling function for positive sets, i.e. MinPooling, with MaxPooling, MeanPooling, SumPooling, GAP (Li et al., 2016), FSP (Zhang et al., 2018), and PSWE (Naderializadeh et al., 2021). The results are denoted in Table LABEL:tab:exp_pooling_results_positive.
@@
MinPooling 0.333$\pm$0.006 0.284$\pm$0.003 0.195$\pm$0.006 0.191$\pm$0.010 0.494$\pm$0.007 0.484$\pm$0.007 0.324$\pm$0.009 0.211$\pm$0.001
MeanPooling 0.335$\pm$0.003 0.284$\pm$0.005 0.197$\pm$0.006 0.194$\pm$0.007 0.492$\pm$0.009 0.481$\pm$0.006 0.324$\pm$0.013 0.213$\pm$0.002
SumPooling 0.336$\pm$0.006 0.286$\pm$0.004 0.195$\pm$0.007 0.186$\pm$0.008 0.493$\pm$0.009 0.481$\pm$0.008 0.320$\pm$0.013 0.212$\pm$0.001
GAP 0.336$\pm$0.006 0.289$\pm$0.005 0.196$\pm$0.006 0.186$\pm$0.005 0.492$\pm$0.010 0.482$\pm$0.007 0.318$\pm$0.009 0.210$\pm$0.002
FSP 0.336$\pm$0.004 0.285$\pm$0.003 0.197$\pm$0.006 0.191$\pm$0.010 0.494$\pm$0.012 0.484$\pm$0.008 0.323$\pm$0.014 0.218$\pm$0.002
PSWE 0.333$\pm$0.007 0.285$\pm$0.005 0.198$\pm$0.005 0.187$\pm$0.006 0.494$\pm$0.009 0.484$\pm$0.005 0.322$\pm$0.008 0.211$\pm$0.002
@@ Appendix E Examples of data augmentation methods
ronaldo poaches the points as owen tastes first real. ronaldo’s scoring prowess was on full display as he secured the victory for real, while owen experienced his first taste of success with the team. the serene sunset over the ocean painted the sky with hues of orange and pink, casting a peaceful aura over the coastal town.
i quite often find that when a film doesn’t evolve around a famous actor or actress but rather a story or style. focusing on a compelling story or unique cinematic style can often lead to a more engaging movie experience, regardless of the star power of the actors involved. the cat jumped over the fence and landed gracefully in the neighbor’s yard.
finally, a way to Google your hard drive You can usually find information more easily on the vast internet. with this new tool, you’ll be able to search your hard drive just as efficiently as you browse the vast expanse of the internet. the sound of laughter echoed through the empty park.
bioshield effort is inadequate, a study says experts say the bioshield program will not be adequate to create new vaccines and drugs to protect the nation against either a biological attack or natural epidemic. the study’s findings cast doubt on the effectiveness of the bioshield program in developing the necessary vaccines and drugs to safeguard the nation from biological threats, according to experts. the chef carefully placed the garnish on the plate, completing the presentation of the dish.
@@
ronaldo poaches the points as owen tastes first real. cristiano cambone this rebounds one bellamy distinctly fourth gives. maradona lem following 18 own ritchie mixes 1998 rather.
i quite often find that when a film doesn’t evolve around a famous actor or actress but rather a story or style. never seemed ones way has he with screenplay reason we alter apart with notable film rather girlfriend only meant with explains rather simple. get rarely perhaps still actually turn put episode come nothing conversely instead put besides scriptwriter there ms. while instance put inspired there quirky.
finally, a way to Google your hard drive You can usually find information more easily on the vast internet. once both with it move online get turning stopping maybe if instead way account still quickly which this wealth websites. turned instead put once rest lets want lot hard lot these less still link longer none last following boasts msn.
bioshield effort is inadequate, a study says experts say the bioshield program will not be adequate to create new vaccines and drugs to protect the nation against either a biological attack or natural epidemic. nobelity helping part sufficiently both with suggests explained researchers know this nobelity planning ready any even providing move rather its doses made treatment move threatened this entire losing however with possessed fire rather unique malaria. circumvents promises once poorly instead put applied nothing impact n’t following circumvents expanded step thought cannot suitable rest core introduced medication few antiviral rest need following strongest twice as put analyzing enemy there hence spreads.
@@
ronaldo poaches the points as owen tastes first real. ronaldinho strafaci which point also keane taste second kind. artd newsfeatures yahd nzse40 404-526-5456 3,399 leel arabised highjump.
i quite often find that when a film doesn’t evolve around a famous actor or actress but rather a story or style. me very sometimes how but then an movie if know evolving where an famed starred instead actor though merely an book instead styles. statistique kunzig -96 deylam suhn micale lacques nzse40 xinglong myoo fdch lacques -5000 basse-normandie zahr debehogne cetacea lacques arabised zahr tih.
finally, a way to google your hard drive you can usually find information more easily on the vast internet. again . an so take yahoo my turn drives ’ll able typically how source than either . which surrounding web. sahk lacques -96 yahd raht 1-732-390-4480 milanka musicnotes.com 1-732-390-4480 photoexpress suhn deylam tahn 100.00 zhahn romanosptimes.com newsfeatures puh condita.
bioshield effort is inadequate, a study says experts say the bioshield program will not be adequate to create new vaccines and drugs to protect the nation against either a biological attack or natural epidemic. toshka efforts this insufficient . an studies thinks specialists believe which toshka programs would be not sufficient take creating for vaccine well drug take protecting which country opponents instead an chemical attacks instead abundant outbreak. chivenor wmorrisglobe.com nehn europe/africa sahk lacques ortsgemeinden telegram.com gahl e-bangla newsfeatures chivenor sholdersptimes.com ortsgemeinden milanka 47-42-80-44 sylviidae yahd troo yanow chron.com bizmags tilove yahd telegram.com newsfeatures juh speech/language tvcolumn lacques fassihi 2378 zahr berris bse-100.
@@ Appendix F Statistical Significance Tests towards Topic Quality
In this appendix, to more rigorously assess the quality of our generated topics compared with baseline models, we conduct a statistical t-test for performance comparison in Section 5. In particular, we select the second-best method in all Tables, i.e. NTM+CL, Linear-$\alpha$ = 0.5, NTM+CL, NTM+CL, GAP, and GAP in Table LABEL:tab:exp_dc, LABEL:tab:exp_ablation_objective_control_strategy, 2, 3, LABEL:tab:exp_pooling_results_negative, and LABEL:tab:exp_pooling_results_positive, respectively. We provide our obtained $p$-values in Table 15, 16, 17, 18, 19, and 20. The results show that our improvement is statistically significant with all $p$-value $<0.05$.